• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 08:10:29 on July18,2003

    Entry number 104416

    keyword=owl shift summary

    owl shift run summary

    0000:0100 Moller measurement = 72%, also see septum heating

    0100 MCC called to inform us that beam was being restored, setting beam position etc.

    0120 beam back, 2 microA, raster off, empty target, preparing to do HARP scan
    0125 software problems see entries: 104369, 104370, 104371
    0227 asked for beam off while trying to fix the medm problem
    0327 problem fixed, 104373
    0353 checked everything, documented everything, reopened all windows, asked for beam
    0415:0513 did HARP scans to check beam position and width: 104378, 104383, 104392, 104397. see also ELOG about aspect ratio
    0534 checked spot size (good), MCC called QE measurement, no beam: 104400
    0540 bleedthrough measurement
    0609 MCC called, RF IOC crashed
    0705 target pol = 41%, see entry 104413
    0748 beam back, MCC working on beam tune